### v2

@@ Abstract
Robots can use Visual Imitation Learning (VIL) to learn everyday tasks from video demonstrations. However, translating visual observations into actionable robot policies is challenging due to the high-dimensional nature of video data. This challenge is further exacerbated by the morphological differences between humans and robots, especially when the video demonstrations feature humans performing tasks. To address these problems we introduce Visual Imitation lEarning with Waypoints (VIEW), an algorithm that significantly enhances the sample efficiency of human-to-robot VIL. VIEW achieves this efficiency using a multi-pronged approach: extracting a condensed prior trajectory that captures the demonstrator’s intent, employing an agent-agnostic reward function for feedback on the robot’s actions, and utilizing an exploration algorithm that efficiently samples around waypoints in the extracted trajectory. VIEW also segments the human trajectory into grasp and task phases to further accelerate learning efficiency. Through comprehensive simulations and real-world experiments, VIEW demonstrates improved performance compared to current state-of-the-art VIL methods. VIEW enables robots to learn a diverse range of manipulation tasks involving multiple objects from arbitrarily long video demonstrations. Additionally, it can learn standard manipulation tasks such as pushing or moving objects from a single video demonstration in under 30 minutes, with fewer than 20 real-world rollouts. Code and videos here: https://collab.me.vt.edu/view/
@@ 1 Introduction
Imagine teaching a person to pick up a cup placed on a table. The quickest method is often to physically demonstrate this task. Through physical demonstration, the observer can discern which object to pick up and how to manipulate that object. Humans efficiently learn everyday tasks in this way, including moving items, pouring tea, or stirring the contents of a pan.
Teaching robots these same tasks, however, proves to be more cumbersome. Typically, robots employ either Imitation Learning (IL) or Reinforcement Learning (RL) methods. IL generally requires many demonstrations from humans to obtain effective policies fang2019survey ; hussein2017imitation . During this process, the human teacher often needs to kinesthetically guide or teleoperate the robot to show it exactly what actions it should take. On the other hand, RL methods require a substantial number of rollouts for robots to perform even simple tasks kober2013reinforcement ; morales2021survey . Additionally, defining appropriate reward functions for reinforcement learning poses a challenge, particularly in unstructured everyday settings dulac2021challenges .
@@
Accordingly, the second part of VIEW focuses on iteratively improving the robot’s prior and correctly completing the task. We develop sampling strategies so that the robot can intelligently explore around its waypoints. This includes waypoints where the robot needs to grasp an object (e.g., pick up the cup) and waypoints where the robot is manipulating that object (e.g., carrying the cup to a goal location). Again, we rely on our hypothesis: instead of reasoning about every aspect of the video, we focus on the object’s location in the waypoint frames. This leads to an iterative learning process where the robot corrects its initial trajectory and eventually completes the original task shown in the video. Overall, VIEW enables robot arms to efficiently learn everyday tasks such as picking, pushing, or moving objects, requiring fewer than $20$ real-world trials and less than $30$ minutes from demonstration collection to successful task execution. Additionally, our method enables robots to learn from long horizon videos that involve a combination of tasks — such as moving a cup and pouring tea into it, or placing multiple objects in a pan — using no additional information besides the initial human demonstration video.
@@ 2 Related Work
Several approaches closely align with our method, focusing on human-to-robot imitation learning lee2022learning ; jain2024vid2robot ; jin2020geometric ; sieb2020graph ; alakuijala2023learning ; sermanet2017unsupervised ; chane2023learning ; Patel2022 ; shaw2024learning ; whirl . These methods extract meaningful representations of a task from videos chane2023learning or use neural networks to learn reward functions from the videos to facilitate reinforcement learning sermanet2017unsupervised ; alakuijala2023learning . Despite the promise shown by many of these methods, they share a common challenge: the necessity for a substantial number of robot rollouts in real-world scenarios to learn tasks effectively. One particularly similar approach here is WHIRL whirl , which mirrors our method but employs a variational autoencoder-based exploration exploitation strategy. As we will show, WHIRL requires a large number of rollouts to converge, and struggles to scale for long horizon tasks. Additionally, it relies on video inpainting lee2019copy that can be time-consuming. Our proposed method VIEW aims to solve these challenges by segmenting the task and sequentially solving it: in our experiments, we will directly compare VIEW and WHIRL in terms of task success and training time.
@@
Human activity recognition. Many of the methods discussed above rely on human intent and activity recognition for enabling robots to understand object affordance. Within the realm of robot manipulation, numerous studies have proposed methods for discerning human actions and activities from video footage eze2024learning ; luo2023learning ; koppula2013learning ; koppula2015anticipating . Some works extend human posture tracking to identify fine-grained activities within limited spatial contexts ma2018region . However, merely detecting human presence is insufficient to learn from human videos. To bridge this gap, using annotated video datasets such as SomethingSomething goyal2017something , YouCook das2013thousand , ActivityNet caba2015activitynet , or the 100 Days of Hands (100DOH) shan2020understanding becomes instrumental. The 100DOH dataset is particularly valuable due to its detailed object interaction annotations. Building upon prior works, our approach utilizes the 100DOH framework to extract crucial data on hand positioning and interactions with objects.
@@ 3 Problem Statement
We consider manipulation tasks in unstructured environments. First a human teacher physically demonstrates their desired task within the robot’s workspace. During this demonstration the robot is moved out of the way (i.e., the human does not interact with the robot) and the robot records the human’s behavior with a stationary RGB-D camera. After the demonstration is complete the video is provided to the robot, and the robot must learn how to replicate the same task based on this single video. We highlight that the human and robot have morphological differences — e.g., the human’s hand is different from the robot’s gripper — and so the way the human performed the task may not transfer directly to the robot arm.
@@
Video Demonstration. The robot learns from a single video demonstration ($V_{i}$) of a human performing the task ($\tau_{i}$) captured using a stationary RGB-D camera. The human does not interact with the robot beyond providing this video. Although the human only provides one video for task $\tau_{i}$, they may provide demonstrations for multiple tasks: e.g., the robot could receive a set of $n$ videos $V_{1},\ldots,V_{n}$ for $n$ different tasks $\tau_{1},\ldots,\tau_{n}$. The robot’s objective is to map each video into a trajectory that completes the demonstrated task.

### v3

@@ Abstract
Robots can use Visual Imitation Learning (VIL) to learn manipulation tasks from video demonstrations. However, translating visual observations into actionable robot policies is challenging due to the high-dimensional nature of video data. This challenge is further exacerbated by the morphological differences between humans and robots, especially when the video demonstrations feature humans performing tasks. To address these problems we introduce Visual Imitation lEarning with Waypoints (VIEW), an algorithm that significantly enhances the sample efficiency of human-to-robot VIL. VIEW achieves this efficiency using a multi-pronged approach: extracting a condensed prior trajectory that captures the demonstrator’s intent, employing an agent-agnostic reward function for feedback on the robot’s actions, and utilizing an exploration algorithm that efficiently samples around waypoints in the extracted trajectory. VIEW also segments the human trajectory into grasp and task phases to further accelerate learning efficiency. Through comprehensive simulations and real-world experiments, VIEW demonstrates improved performance compared to current state-of-the-art VIL methods. VIEW enables robots to learn manipulation tasks involving multiple objects from arbitrarily long video demonstrations. Additionally, it can learn standard manipulation tasks such as pushing or moving objects from a single video demonstration in under 30 minutes, with fewer than 20 real-world rollouts. Code and videos here: https://collab.me.vt.edu/view/
@@ 1 Introduction
Imagine teaching a person to pick up a cup placed on a table. The quickest method is often to physically demonstrate this task. Through physical demonstration, the observer can discern which object to pick up and how to manipulate that object. Humans efficiently learn everyday tasks in this way, including moving items, making tea, or stirring the contents of a pan.
Teaching robots these same tasks, however, proves to be more cumbersome. Typically, robots employ either Imitation Learning (IL) or Reinforcement Learning (RL) methods. IL generally requires many demonstrations from humans to obtain effective policies fang2019survey ; hussein2017imitation . During this process, the human teacher often needs to kinesthetically guide or teleoperate the robot to show it exactly what actions it should take. On the other hand, RL methods require a substantial number of rollouts for robots to perform even simple tasks kober2013reinforcement ; morales2021survey . Additionally, defining appropriate reward functions for reinforcement learning poses a challenge, particularly in unstructured environments dulac2021challenges .
@@
Accordingly, the second part of VIEW focuses on iteratively improving the robot’s prior and correctly completing the task. We develop sampling strategies so that the robot can intelligently explore around its waypoints. This includes waypoints where the robot needs to grasp an object (e.g., pick up the cup) and waypoints where the robot is manipulating that object (e.g., carrying the cup to a goal location). Again, we rely on our hypothesis: instead of reasoning about every aspect of the video, we focus on the object’s location in the waypoint frames. This leads to an iterative learning process where the robot corrects its initial trajectory and eventually completes the original task shown in the video. Overall, VIEW enables robot arms to efficiently learn manipulation tasks such as picking, pushing, or moving objects, requiring fewer than $20$ real-world trials and less than $30$ minutes from demonstration collection to successful task execution. Additionally, our method enables robots to learn from long horizon videos that involve a combination of tasks — such as moving a cup and dispensing tea into it, or placing multiple objects in a pan — using no additional information besides the initial human demonstration video.
@@ 2 Related Work
Several approaches closely align with our method, focusing on human-to-robot imitation learning lee2022learning ; jain2024vid2robot ; jin2020geometric ; sieb2020graph ; alakuijala2023learning ; sermanet2017unsupervised ; chane2023learning ; Patel2022 ; shaw2024learning ; whirl . These methods extract meaningful representations of a task from videos chane2023learning or use neural networks to learn reward functions from the videos to facilitate reinforcement learning sermanet2017unsupervised ; alakuijala2023learning . Despite the promise shown by many of these methods, they share a common challenge: the necessity for a substantial number of robot rollouts in real-world scenarios to learn tasks effectively. One particularly similar approach here is WHIRL whirl , which mirrors our method but employs a variational autoencoder-based exploration exploitation strategy. As we will show, WHIRL requires a large number of rollouts to converge, and struggles to scale for long horizon tasks. Additionally, this approach also requires video inpainting lee2019copy , where the human must be removed from each frame in the demonstration video, and the robot must be removed from each rollout frame before reward computation. Given that inpainting is highly GPU-intensive, this process can require powerful GPUs and lead to substantial wait times between rollouts. Our proposed method, VIEW, addresses these challenges by segmenting and sequentially solving the task, reducing computational demands. In our experiments, we will directly compare VIEW and WHIRL in terms of task success and training time to highlight these improvements.
@@
Human activity recognition. Many of the methods discussed above rely on human intent and activity recognition for enabling robots to understand object affordance. Within the realm of robot manipulation, numerous studies have proposed methods that use annotated video datasets such as SomethingSomething goyal2017something , YouCook das2013thousand , ActivityNet caba2015activitynet , or the 100 Days of Hands (100DOH) shan2020understanding . The 100DOH dataset is particularly valuable due to its detailed object interaction annotations. Building upon prior works, our approach utilizes the 100DOH framework to extract crucial data on hand positioning and interactions with objects.
@@ 3 Problem Statement
We consider single object manipulation tasks in unstructured environments. First a human teacher physically demonstrates their desired task within the robot’s workspace. During this demonstration the robot is moved out of the way (i.e., the human does not interact with the robot) and the robot records the human’s behavior with a stationary RGB-D camera. After the demonstration is complete the video is provided to the robot, and the robot must learn how to replicate the same task based on this single video. We highlight that the human and robot have morphological differences — e.g., the human’s hand is different from the robot’s gripper — and so the way the human performed the task may not transfer directly to the robot arm.
@@
Video Demonstration. The robot learns from a single video demonstration ($V_{i}$) of a human performing the task ($\tau_{i}$). This video is captured using a stationary RGB-D camera that is set up such that the human and the object they manipulate are visible at all times. The human does not interact with the robot beyond providing this video. Although the human only provides one video for task $\tau_{i}$, they may provide demonstrations for multiple tasks: e.g., the robot could receive a set of $n$ videos $V_{1},\ldots,V_{n}$ for $n$ different tasks $\tau_{1},\ldots,\tau_{n}$. The robot’s objective is to map each video into a trajectory that completes the demonstrated task.
Tasks. Our focus is primarily on tasks that involve manipulating a single object at a time. We consider tasks such as picking, pushing, or moving an object within the robot’s workspace. While multi-object tasks are allowed, they must be decomposed into multiple sequential single-object manipulation tasks, similar to those mentioned above.